OpenStudy (sonyalee77):

Mary’s yard is a mess. She needs to hire someone to prune her trees and shrubs. A landscaping service she calls quotes her a price of $15 consultation fee plus $8 an hour for the actual work. Mary’s neighbor has offered to help her out. She doesn’t charge a consultation fee, but does charge $10 an hour for her work. Who do you think Mary should hire for a 3-hour job? When, if at all, would it be more economical to hire the other service. (Set up and solve a system of equations to answer this question.)

OpenStudy (sonyalee77):

a. Mary should hire her neighbor; if there was more than 7.5 hours of work she should hire the landscaping service. b. Mary should hire her neighbor; if there was more than 4.2 hours of work she should hire the landscaping service. c. Mary should hire the landscaping service; if there was more than 7.5 hours of work she should hire her neighbor. d. Mary should hire the landscaping service; if there was more than 4.2 hours of work she should hire her neighbor.

OpenStudy (sonyalee77):

@texaschic101 @Zale101

OpenStudy (sonyalee77):

@helpme1.2 @jigglypuff314

OpenStudy (sonyalee77):

@ganeshie8 @amistre64

OpenStudy (sonyalee77):

pleasse help

OpenStudy (sonyalee77):

@anonymous97

OpenStudy (anonymous):

let x = hours of work first question: who costs less for x = 3 second one: after how many hours do their services end up costing the same? S = service = 15 + 8x N = neighbor = 0 + 10x = 10x 1) for x = 3 which is better? 2) what is x when: 15 + 8x = 10x ?

OpenStudy (emilylauren):

The answer is A I will explain if you'd like

OpenStudy (anonymous):

i agree that it is A

OpenStudy (emilylauren):

Euler explained it well, make sure you understand how to do the problem :)

OpenStudy (sonyalee77):

the answer is 7.5?

OpenStudy (sonyalee77):

thank you very much! :D

OpenStudy (emilylauren):

7.5 hours is when it becomes more cost effective to hire the landscaping service than the neighbor. So, yes.

OpenStudy (sonyalee77):

Emily Lauren I would fan you but my computer wont show the fan button, also cant see the smart scores :( .... sorry i'll keep trying though.